OpenStudy (anonymous):

what are real solution of the equation |x|^2+2|x|-3=0

OpenStudy (anonymous):

graph it

OpenStudy (anonymous):

or if you want to do it by analysis, split the absolute value up into two cases where x is negative and where x is positive.

OpenStudy (anonymous):

already did it

OpenStudy (agent0smith):

Hint: you can drop the absolute value signs on the x^2.. squaring a number makes it positive anyway

OpenStudy (kainui):

Solve this for |x| just like you would for x normally. Don't let the little bars fool you, it's almost identical to any other way, use the quadratic formula or complete the square.

OpenStudy (kainui):

Dropping the absolute value signs here isn't going to help you.

OpenStudy (agent0smith):

you don't need to anyway just factor it |x|^2+2|x|-3=0 (|x| + 3) (|x| -1) = 0

OpenStudy (kainui):

That's what I'm saying... You're just repeating me and giving useless hints to boot.

OpenStudy (agent0smith):

|x| + 3 = 0 |x| -1=0

OpenStudy (agent0smith):

"use the quadratic formula or complete the square." yes i repeated you exactly.

OpenStudy (kainui):

Good job you solved a girl's quadratic formula for her without her doing any work.

OpenStudy (anonymous):

got it

OpenStudy (agent0smith):

x^2+2x-3=0 would factor as (x + 3) (x -1) = 0 right?

OpenStudy (agent0smith):

|x| + 3 = 0 |x| -1=0 keep in mind only one of those has a solution...

OpenStudy (anonymous):

so we are going to solve it with out removing absolute values

OpenStudy (agent0smith):

well with the two above you need to remove them...

OpenStudy (agent0smith):

well one of them, hopefully you'll notice which one has a solution and which doesn't

OpenStudy (anonymous):

+-1

OpenStudy (agent0smith):

Good :)

OpenStudy (anonymous):

can syou show me a complete solution of the equation i want to confirm

OpenStudy (agent0smith):

|x|^2+2|x|-3=0 factor it first... as above (|x| + 3) (|x| -1) = 0 |x| + 3 = 0 ... |x| = -3 no solution |x| -1=0 ... |x| = 1 ... x = 1 or -1

OpenStudy (agent0smith):

Post a new question.

OpenStudy (anonymous):

http://prntscr.com/276aix

OpenStudy (anonymous):

can you help with this one i actually measure angle x instead of solving it

ganeshie8 (ganeshie8):

find interior angle of pentagon find exterior angle of pentagon

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

u should be knowing those two angles right ? :)

ganeshie8 (ganeshie8):

exterior angle of pentagon = 360/5 = 72

ganeshie8 (ganeshie8):